OpenStudy (anonymous):

Find the exact value of sin (arccos (3/5)). For full credit, explain your reasoning.

myininaya (myininaya):

So if we have x=arccos(3/5) then cos(x)=3/5. Draw a right triangle and use cos(x)=3/5 to find the other missing leg. then evaluate sin(x)

OpenStudy (anonymous):

Myininaya has it down, abc. You should follow myininaya's hint.

OpenStudy (anonymous):

im sorry i still dont understand it could you explain it more? like is it like pythagorean therom?

OpenStudy (anonymous):

Exactly. recall that cosx is just adjacent over the hypotenuse. So, your adj side is 3 and your hypotenuse is 5. Now just find the missing side by applying pythagorean's theorem and then go back and answer your question.

OpenStudy (anonymous):

would it be 4/5?

OpenStudy (phi):

yes For full credit, explain your reasoning.

OpenStudy (anonymous):

I drew a right triangle I used cos(x)=3/5 to find the other missing leg. I know that cosx is just adjacent over the hypotenuse. So, the adj side is 3 and your hypotenuse is 5. I found the missing side by applying pythagorean's theorem and my answer is 4/5

OpenStudy (anonymous):

@phi is this a good answer?

OpenStudy (phi):

yes, sounds good (but I would leave out the "just" in "just adjacent over hypotenuse" I found the missing side by applying pythagorean's theorem I would add "and it is 4" then add the definition of the sin: opposite over adjacent

OpenStudy (anonymous):

okay thank you so much everyone for all of your help!!!